A sample of thallium-208 has a half-life of 3.053 minutes. What fraction of the parent material will remain after 12.212 minutes?


1
Expert's answer
2021-01-12T12:11:19-0500

"N=N_0(\frac{1}{2})^{\frac{t}{t_{1/2}}}" , where N is the amount at given time (in this case - the fraction), N0 is the initial amount (in this case it is just equal to 1 because we are asked to find the fraction only), t is time, and t1/2 is the half-life. Therefore,


"N=1\times(\frac{1}{2})^{\frac{12.212min}{3.053min}}=0.0625"


Answer: 0.0625